# Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game

### v5

@@ Appendix B Disagreements in Annotations
There was certain extent of disagreements between Semantic and Encyclopedic Knowledge. For instance [’PIKE’, ’SPLIT’, ’STRADDLE’, ’TUCK’] are GYMNASTICS POSITIONS and requires domain specific knowledge so could be thought of as Encyclopedic knowledge but it could be classified under Semantic Knowledge (Type Of relation) as many of these words appear in Wordnet. Some disagreements also occurred between between Associative and Encyclopedic Knowledge. For instance here the shared property for [’BASE’, ’BOND’, ’ELEMENT’, ’SOLUTION’] being CHEMISTRY TERMS requires using Associative Knowledge but this could still require Encyclopedic Knowledge about Chemistry. However since we consider Encyclopedic knowledge only as ones related to a Knowledge Base and entities instead of domain concepts we treat this as Associative Knowledge. Finally due to their colloquial use in the English language sometimes there can be confusion amongst what Semantic and Associative Knowledge. For instance [’BOMB’, ’DUD’, ’FLOP’, ’LEMON’] can be thought of synonyms of FAILURE and hence fall under category of Semantic Knowledge , however Lemon is rarely used for Failure and requires using connotative knowledge (shared property) and hence falls more appropriately under Associative Knowledge.
@@ Appendix C Red Herrings
In the puzzle in Figure 6, a red herring category is present. In the highest performing models Claude 3 Opus and GPT-4o created a category called "Milk" with Whole, Skim, and Soy and included a random fourth word that did not fit. Each of these three words, however, belongs to a different category: Whole to Kinds of Numbers, Skim to Touch Lightly, and Soy to Sauces in Chinese Cuisine. In other puzzles including a red herring category like this one, all models make similar rationalizations.
The game in Figure 7 is an example of a game with a red herring word. The five words that appear as though they belong together are outlined in red. However, Mistletoe, Reindeer, Snowman, and Stocking form the Christmas Related category, while Candy Cane belongs to the category Things with Stripes. In this game, GPT-4o, Gemini 1.5 Pro, and Llama 3 70B all made the mistake of grouping Candy Cane with some combination of three of the other Christmas-related words.
@@ D.1 Humans
The frequency of clustering scores for novice human players in 100 games and scores for GPT-4o in the same games are shown in Table 4. The frequency of clustering scores for expert human players in 50 games and scores for GPT-4o in the same games are in Table 5.
Figures 8 and 9 show the distribution for the weighted clustering scores of GPT-4o against novice humans and expert humans, respectively. Finally, Figures 10 and 11 depict the performance of novice and expert humans, respectively, by reasoning type from our taxonomy of knowledge. Because the total counts of types of reasoning required across the 400 (novice) and 200 (expert) categories are unbalanced, the count of the categories reasoned correctly is shown above each bar.
GPT-4o | Novice Humans | |||
0 | 31 | 30 | |||
1 | 35 | 39 | |||
2 | 29 | 13 | |||
3 | 2 | 0 | |||
4 | 8 | 18 |
GPT-4o | Expert Humans | |||
0 | 16 | 2 | |||
1 | 18 | 7 | |||
2 | 10 | 9 | |||
3 | 1 | 0 | |||
4 | 5 | 32 |
@@ D.2 LLMs
Table 6 shows the frequency of the unweighted clustering scores (number of categories correctly grouped) for each LLM. The total number of games played by each model is 200. Table 7 is slightly different and shows the frequency of categorical reasoning scores (the categories correctly grouped and reasoned) for each model. Because a caveat for receiving a categorical reasoning score greater than 0 is matching gold categories, a score of 0 is more common than in the unweighted clustering scores.
Figure 12 shows the performance of each LLM by reasoning type from our taxonomy of knowledge. Because the total counts of types of reasoning required across the 800 categories are unbalanced, the count of the categories reasoned correctly is shown above each bar.
Gemini 1.5 Pro | Claude 3 Opus | GPT-4o | Llama 3 70B | |||
0 | 100 | 68 | 53 | 84 | |||
1 | 64 | 72 | 77 | 73 | |||
2 | 26 | 37 | 48 | 32 | |||
3 | 1 | 10 | 6 | 5 | |||
4 | 9 | 13 | 16 | 6 |
Gemini 1.5 Pro | Claude 3 Opus | GPT-4o | Llama 3 70B | ||
0 | 116 | 80 | 59 | 98 | ||
1 | 59 | 66 | 87 | 73 | ||
2 | 18 | 36 | 41 | 23 | ||
3 | 7 | 13 | 11 | 4 | ||
4 | 0 | 5 | 2 | 2 |
@@ Appendix E Human Evaluation Interface
Figure 13 shows the two main screens of the evaluation interface provided to both novice and expert human evaluators. (a) is the instruction screen, while (b) is an example of a game screen after the user hits the "Play" button. To solve the game in one shot, all 16 words from a game are displayed on the screen in separate boxes, with one drop-down per box. The drop-down consists of four labels: Group 1, Group 2, Group 3, and Group 4. The user’s job is to create 4 groups of 4 words using the given labels. Because the groups are chosen from a drop-down menu where the default option is Group 1, a clustering score of 3 is impossible.

### v6

@@ Appendix B Disagreements in Annotations
There was some disagreement between Semantic and Encyclopedic Knowledge labeling. One example is the grouping pike, split, straddle, tuck which are “Gymnastics Positions" and requires domain-specific knowledge, meaning it could be thought of as Encyclopedic. However, it could also be classified under Semantic Relations (Type Of relation) as many of these words appear in WordNet. There can be confusion about what is considered to be Semantic or Associative Knowledge. For example, card, chocolate, heart, rose might be thought of to be semantically related as gifts given to a loved one and hence labeled under semantic relations; however, they can be viewed as elements of specific situations or thematic context such as “Seen on Valentine’s Day". Hence it falls more appropriately under Associative Relations.
@@ Appendix C Red Herrings
In the puzzle in Figure 8, a red herring category is present. Gemini 1.5 Pro created a category called "Milk" with Whole, Skim, and Soy and included a random fourth word that did not fit. Each of these three words, however, belongs to a different category: Whole to Kinds of Numbers, Skim to Touch Lightly, and Soy to Sauces in Chinese Cuisine. In other puzzles including a red herring category like this one, most models make similar rationalizations.
The game in Figure 9 is an example of a game with a red herring word. The five words that appear as though they belong together are outlined in red. However, Mistletoe, Reindeer, Snowman, and Stocking form the “Christmas Related" category, while Candy Cane belongs to the category “Things with Stripes". In this game, all models except Claude 3.5 Sonnet made the mistake of grouping Candy Cane with some combination of three of the other Christmas-related words.
@@ D.1 Humans
The frequency of clustering scores for novice human players in 100 games and scores for Claude 3.5 Sonnet in the same games are shown in Table 3. The frequency of clustering scores for expert human players in 50 games and scores for Claude 3.5 Sonnet in the same games are in Table 4.
Figures 10 and 11 show the distribution of the weighted clustering scores for Claude 3.5 Sonnet against novice humans and expert humans, respectively.
Novice Humans | |||||
| 0 | 17 | 30 | |||||
| 1 | 41 | 39 | |||||
| 2 | 27 | 13 | |||||
| 3 | 3 | 0 | |||||
| 4 | 12 | 18 |
Expert Humans | |||||
| 0 | 6 | 2 | |||||
| 1 | 20 | 7 | |||||
| 2 | 13 | 9 | |||||
| 3 | 1 | 0 | |||||
| 4 | 10 | 32 |
Gemini 1.5 Pro | Claude 3.5 Sonnet | GPT-4o | Llama 3.1 405B | Mistral Large 2 | |||
| 0 | 161 | 87 | 112 | 146 | 185 | |||
| 1 | 153 | 138 | 140 | 150 | 149 | |||
| 2 | 90 | 117 | 111 | 93 | 80 | |||
| 3 | 10 | 17 | 10 | 2 | 3 | |||
| 4 | 24 | 79 | 65 | 47 | 21 |
Gemini 1.5 Pro | Claude 3.5 Sonnet | GPT-4o | Llama 3.1 405B | Mistral Large 2 | |||
| 0 | 203 | 101 | 131 | 176 | 215 | |||
| 1 | 154 | 152 | 157 | 151 | 144 | |||
| 2 | 59 | 109 | 94 | 73 | 61 | |||
| 3 | 16 | 53 | 36 | 25 | 13 | |||
| 4 | 6 | 23 | 20 | 13 | 5 |
@@ D.2 LLMs
Table 5 shows the frequency of the unweighted clustering scores (number of categories correctly grouped) for each LLM. The total number of games played by each model is 438. Table 6 is slightly different and shows the frequency of categorical reasoning scores (the categories correctly grouped and reasoned) for each model. Because a caveat for receiving a categorical reasoning score greater than 0 is matching gold category words and names, a score of 0 is more common than in the unweighted clustering scores.
@@ Appendix E Human Evaluation Interface
Figure 12 shows the two main screens of the evaluation interface provided to both novice and expert human players. (a) is the instruction screen, while (b) is an example of a game screen after the user hits the "Play" button. To solve the game in one shot, all 16 words from a game are displayed on the screen in separate boxes, with one drop-down per box. The drop-down consists of four labels: Group 1, Group 2, Group 3, and Group 4. The user’s job is to create 4 groups of 4 words using the given labels. Because the groups are chosen from a drop-down menu where the default option is Group 1, a clustering score of 3 is impossible.